OpenStudy (anonymous):

graph using the slope and the y intercept x+5y=20

OpenStudy (anonymous):

x + 5y = 20 5y = -x + 20 y = -x/5 + 4 so slope = -1/5

OpenStudy (lalaly):

slope is -1/5 and the y intercept is 4...its a straight line

OpenStudy (asadkarim7):

y = -1/5 x + 4 m= -1/5 intercept = 4

OpenStudy (anonymous):

now put x = 0 y = 0 + 4 y = 4 is the y intercept

OpenStudy (anonymous):

thanx yall are my hero!!!!!!!!!!!medals for alll

OpenStudy (anonymous):

hey dinahr, thx for the medal..... can u become my fan too pls !!

OpenStudy (anonymous):

thanks for becoming my fan Asad..... {{ :})

OpenStudy (anonymous):

no prob...im the #1 fan...lol

OpenStudy (anonymous):

thx , just click the fan button pls....

OpenStudy (anonymous):

k in a min

OpenStudy (anonymous):

imma fan now!!!!

OpenStudy (anonymous):

thanks dinahnr